OpenStudy (anonymous):

Totally confused!! Simplify the trigonometric expression. sec Θ cos Θ Answer choices: a. tan Θ b. 1 c. cot Θ d. sin Θ

OpenStudy (anonymous):

What is another way to represent secant? :)

OpenStudy (whpalmer4):

Do you know the definition of \(\sec \theta\)?

OpenStudy (anonymous):

Not really. I am very new at this and have no idea what I'm doing.

OpenStudy (anonymous):

It's beginner's Trigonometry. What's another way to represent Secant? (As a fraction)

OpenStudy (whpalmer4):

Okay, well, time to start learning :-) \[\sec \theta = \frac{1}{\cos \theta}\]

OpenStudy (anonymous):

Going off of @whpalmer4 's comment, we have \[1/\cos \theta * \cos \theta = \cos \theta/\cos \theta\] which is what?

OpenStudy (anonymous):

1?

OpenStudy (whpalmer4):

Yes, that's correct.